OpenStudy (sharm_mar31):

how can i factor 125-y cube

OpenStudy (anonymous):

(125-y)^3 = (125-y)(125-y)(125-y). Its already factored to its simplest form Unless they want you to distribute the cubic aspect.

OpenStudy (sharm_mar31):

just like this 125-y^3

OpenStudy (anonymous):

it's like 5^3 - y^3 use a^3 - b^3

OpenStudy (anonymous):

a3 − b3 = (a−b)(a2 + ab + b2)

OpenStudy (anonymous):

an - bn = (a-b)(an-1 + an-2b + an-3b2 + +bn-1)

OpenStudy (sharm_mar31):

thanks!! it really works